OpenStudy (nthenic_oftime):

Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 medals :) Give your answers in exact form. Show your work.

OpenStudy (kkutie7):

correct me if I'm wrong but I'm pretty sure you just solve for x

OpenStudy (pawanyadav):

I think you want the difference of its two roots

OpenStudy (nthenic_oftime):

i didnt think it was JUST for x was it?

OpenStudy (nthenic_oftime):

sorry if this is stupid its SO late here

OpenStudy (kkutie7):

I was assuming the whole y= function (x) thing so y= 0 then solve for possible x int. No? It's late here too I could be wrong

OpenStudy (pawanyadav):

Is it √(2/3)

OpenStudy (kkutie7):

\[x=\frac{-b \pm\sqrt{b^{2}-4ac}}{2a}\] you're missing something

OpenStudy (nthenic_oftime):

@Pawanyadav can you catch me up with howyuou got that ?

OpenStudy (pawanyadav):

I have difference of two roots Let them a,b are two roots Then (a-b)^2 =(a+b)^2 -4ab Here we know sum and product of roots so we can easily find the difference

OpenStudy (kkutie7):

yeah I don't know http://www.wolframalpha.com/input/?i=x+intercepts+of+6x%5E(2)%2B12x%2B5%3D0

Directrix (directrix):

We are looking for the two values of x at which 6x2 + 12x + 5 will equal 0. 6x^2 + 12x + 5 = 0 The Quadratic Formula is needed to find them.

Directrix (directrix):

@nthenic_oftime Do you know how to get started on this problem using the Quadratic Formula?